Belden, Inc. acquires 30 percent of the outstanding voting shares of Sheffield, Inc. on January 1, 2017, for $324,000, which gives Belden the ability to significantly influence Sheffield. Sheffield has a net book value of $792,000 at January 1, 2017. Sheffield's asset and liability accounts showed carrying amounts considered equal to fair values except for a copyright whose value accounted for Belden's excess cost over book value in its 30 percent purchase. The copyright had a remaining life of 16 years at January 1, 2017. No goodwill resulted from Belden's share purchase. Sheffield reported net income of $174,000 in 2017 and $240,000 of net income during 2018. Dividends of $80,000 and $86,000 are declared and paid in 2017 and 2018, respectively. Belden uses the equity method. On its 2018 comparative income statements, how much income would Belden report for 2017 and 2018 in connection with the company's investment in Sheffield? If Belden sells its entire investment in Sheffield on January 1, 2019, for $422,000 cash, what is the impact on Belden's income? Assume that Belden sells inventory to Sheffield during 2017 and 2018 as follows. What amount of equity income should Belden recognize for the year 2018? Year Cost to Belden Price to Sheffield Year-End Balance (at Transfer Price) 2017 $25,760 $46,000 $20,000 (sold in following year) 2018 31,110 61,000 38,000 (sold in following year)
Answer #1

Answer Page No 0 o beguiling income in 2018 Annual amortization - Purchase price - Net booke value Remaining life of copyrighPage No @ Equity income -2018 = 2018 basic equilig income accrual -- 2018 excess fair over book value amortization = 6*240,00. Page No ③ Investment % sheffield - 12/31/18 = 324000 + 46800 -- 24000 + 66,600 – 25800 = $ 387600 Sales price = $422000 Gai1 Page No 4 Beidens ownership = 30% Inbia entily grons profit recognized in 2018 = 304. X8800 $ 2640 2018 - Ending inventor